# Multiple exponential sums and their applications to quadratic congruences

### v3

@@ 5.3. Error estimation
$U$ $=\frac{N^{n}}{p^{m(n+1)}}\sum_{r=0}^{m-2}\sum\limits_{\begin{subarray}{c}(l_{1},..,l_{n})\in\mathbb{Z}^{n}\\ (l_{1},...,l_{n})\not\equiv(0,...,0)\bmod{p}\end{subarray}}\prod_{i=1}^{n}\hat{\Phi}\left(\frac{l_{i}N}{p^{m-r}}\right)e_{p^{m}}\left(-\sum_{i=1}^{n}l_{i}p^{r}x_{0,i}\right)\sum_{\begin{subarray}{c}l=1\\ (l,p)=1\end{subarray}}^{p^{m-r}}E(l_{1}p^{r},...,l_{n}p^{r};p^{m})+O_{\varepsilon}(1)$
(5.8) $\ll\frac{N^{n}}{p^{m(n+1)}}\sum_{r=0}^{m-2}p^{(m+r)n/2+(m-r)}\sum\limits_{\begin{subarray}{c}(l_{1},..,l_{n})\in\mathbb{Z}^{n}\\ (l_{1},...,l_{n})\not\equiv(0,...,0)\bmod{p}\\ Q^{\prime}(L)\equiv 0\bmod p^{m-r-2}\end{subarray}}\prod_{i=1}^{n}\hat{\Phi}\left(\frac{l_{i}N}{p^{m-r}}\right)+O_{\varepsilon}(1).$
@@
$\begin{split}U\ll&\frac{N^{n}}{p^{m(n+1)}}\sum\limits_{r=0}^{m-2}p^{n(m+r)/2}p^{m-r}\sum\limits_{\begin{subarray}{c}(l_{1},..,l_{n})\in\mathbb{Z}^{n}\\ (l_{1},...,l_{n})\not\equiv(0,...,0)\bmod{p}\\ |l_{1}|,...,|l_{n}|\leq L_{r}\\ Q^{\prime}(L)\equiv 0\bmod p^{m-r-2}\end{subarray}}1+O_{\varepsilon}(1).\end{split}$
Let $Q^{\prime}(L)=L^{t}\text{adj}(A)^{t}L=\sum_{i,j}b_{i,j}l_{i}l_{j}$, where the $b_{ij}$’s are only depending on $a_{ij}$’s. Consider $M=\text{max}\{|b_{i,j}|,1\leq i,j\leq n\}$. Now if $L_{r}<\sqrt{p^{m-r-2}/(Mn^{2})}$, then the congruence above can be replaced by the equation $Q^{\prime}(L)=0$. Certainly, this is the case if $N\geq p^{m/2+4m\varepsilon}$ and $m$ is large enough. Hence in this case, we have

### v5

@@ 5.3. Error estimation
$U$ $=\frac{N^{n}}{p^{m(n+1)}}\sum_{r=0}^{m-2}\sum\limits_{\begin{subarray}{c}(l_{1},..,l_{n})\in\mathbb{Z}^{n}\\ (l_{1},...,l_{n})\not\equiv(0,...,0)\bmod{p}\end{subarray}}\prod_{i=1}^{n}\hat{\Phi}\left(\frac{l_{i}N}{p^{m-r}}\right)e_{p^{m}}\left(-\sum_{i=1}^{n}l_{i}p^{r}x_{0,i}\right)\sum_{\begin{subarray}{c}l=1\\ (l,p)=1\end{subarray}}^{p^{m-r}}E(l_{1}p^{r},...,l_{n}p^{r},lp^{r};p^{m})+O_{\varepsilon}(1)$
(5.8) $\ll\frac{N^{n}}{p^{m(n+1)}}\sum_{r=0}^{m-2}p^{(m+r)n/2+(m-r)}\sum\limits_{\begin{subarray}{c}(l_{1},..,l_{n})\in\mathbb{Z}^{n}\\ (l_{1},...,l_{n})\not\equiv(0,...,0)\bmod{p}\\ Q^{\prime}(L)\equiv 0\bmod p^{m-r-1}\end{subarray}}\prod_{i=1}^{n}\hat{\Phi}\left(\frac{l_{i}N}{p^{m-r}}\right)+O_{\varepsilon}(1).$
@@
$\begin{split}U\ll&\frac{N^{n}}{p^{m(n+1)}}\sum\limits_{r=0}^{m-2}p^{n(m+r)/2}p^{m-r}\sum\limits_{\begin{subarray}{c}(l_{1},..,l_{n})\in\mathbb{Z}^{n}\\ (l_{1},...,l_{n})\not\equiv(0,...,0)\bmod{p}\\ |l_{1}|,...,|l_{n}|\leq L_{r}\\ Q^{\prime}(L)\equiv 0\bmod p^{m-r-1}\end{subarray}}1+O_{\varepsilon}(1).\end{split}$
Let $Q^{\prime}(L)=L^{t}\text{adj}(A)^{t}L=\sum_{i,j}b_{i,j}l_{i}l_{j}$, where the $b_{ij}$’s are only depending on $a_{ij}$’s. Consider $M=\text{max}\{|b_{i,j}|,1\leq i,j\leq n\}$. Now if $L_{r}<\sqrt{p^{m-r-1}/(Mn^{2})}$, then the congruence above can be replaced by the equation $Q^{\prime}(L)=0$. Certainly, this is the case if $N\geq p^{m/2+4m\varepsilon}$ and $m$ is large enough. Hence in this case, we have